a stone falls from a ledge and takes 8 seconds to hit the ground.thestone has an original velocity of 0 m/s.how tall is the ledge.: a stone falls from a ledge and takes 8 seconds to hit the ground.thestone has an original velocity of 0 m/s.how tall is the ledge.
Answer by scott8148(2761)

d=(1/2)at^2 __ d=(1/2)(9.8m/sec^2)(8sec)^2 __ d=313.6m